x- 2y = 1
y = -3x + 10
I'm wondering how to solve this linear equation?

Answer by checkley77(12844)

There are 2 ways to solve these equations.
Algebraically & graphically.
x- 2y = 1 or x=2y+1
y = -3x + 10 replace x in this eqaution by (2y+1) & solve for y
y=-3(2y+1)+10
y=-6y-3+10
y+6y=7
7y=7
y=7/7
y=1 ans.
x-2*1=1
x-2=1
x=2+1
x=3 ans.